C- tCu) 4 Ve ) 16-7how did the tutor arrive at 0.2686??

the supporting information: 2$ par, cost 500,000,000, trading for 2.20, dividend of 20cents, growth rate 4%
Answer #1

The Dividend to be included of franking cost=Dividend*(1-franking cost)/(1-tax rate)=0.20*(1-8%)/(1-30%)=0.2686